# Efficient Forkless Blockchain Databases

### v1

@@ 1. Introduction
However, forkless blockchains continue to use legacy databases that were initially designed for forking blockchains only. Recent profiling experiences (Baird et al., 2019) report that transaction execution in forkless blockchains is now primarily dominated by storage access, which consumes up to 75% of the block processing. Moreover, they require nodes to store terabytes of data [^1] Storing the entire history of the Ethereum mainnet currently requires 9 TB of disk space.
@@
Note that the highlighted path for a forkless blockchain is redundant, as it is not needed for the new root $x_{2}$. Retaining the data causes bloat and thus will eventually be pruned (Ethereum, 2025), which typically requires putting the node offline or using costly online algorithms. Moreover, as shown by the dotted arrows, numerous copies of keys along the path need to be copied to create a path from $x_{2}$ to the new value $v^{\prime}_{4}$.
Overall, considering an MPT typically holds hundreds of millions of values, the highlighted inefficiencies lead to considerable computational overhead and redundant bloat if previous versions are not needed.
@@
In this paper, we argue that forkless blockchains need forkless blockchain databases. We thus present a new state database design for high throughput and low storage costs on forkless blockchains. In our approach, we first specialize the state database according to its intended role in the blockchain. The LiveDB is a database for roles such as validator and observer, which require only the state for the latest block on the blockchain. The ArchiveDB is a database for archival nodes that keep all historical states for previous blocks.
The key to our database design is that the LiveDB uses a mutable data structure that allows data to be overwritten, reflecting the constrained versioning occurring in forkless blockchains. Overwriting allows us to reflect the state of the most recent execution without keeping stale states in the database (right MPT in Figure 1). Thus, we perform pruning at no additional cost with LiveDB. We refer to this method as intrinsic pruning that is inherently performed as a consequence of the chosen data structure. Moreover, the LiveDB stores key-value maps as densely as possible. A dense representation of the maps results in smaller disk footprints and faster accesses. This design is in stark contrast to the MPTs that often use pre-existing database solutions such as key-value stores (e.g., LevelDB (Inc., 2023a), RocksDB (Inc., 2023b)) and SQL databases.
@@
LiveDB and ArchiveDB have been implemented in the Fantom Blockchain. Our evaluation shows that LiveDB reduces the storage footprint [^2] Note that even smaller storage decreases imply significant storage reductions linear in the number of nodes in the network because each node replicates the data in a classical L1 blockchain. on the Fantom blockchain by approximately $100\times$ while increasing throughput $10\times$ compared to using the MPT-based Ethereum geth client (Ethereum, 2023). Experiments have been conducted for the public testnet.
Our contributions are summarized as follows:
@@ 2. StateDB Specialization
The StateDB is necessary to verify the validity of transactions, including those that have been skipped due to insufficient funds, incompatible nonces, and other reasons. Since the validator is tasked only with producing the next block, it does not need to retain any historical state; it only requires the latest state of the StateDB. Therefore, after each processed block of transactions, the state of the StateDB can evolve to the latest state. The observer role is similar to the validator role, but it cannot insert new transactions into the ledger; instead, it participates in the consensus protocol and updates its state accordingly.
The archive node role is an extended observer role that keeps the historical states of the blockchain, which can be queried via Remote Procedure Call (RPC). Examples of historical queries include querying an account balance at a specific block height or executing a smart contract at a given block height that queries the information only (but does not change the state). For this reason, it needs to keep the linear evolution of states in the StateDB.
@@
To overcome the shortcomings of this one-size-fits-all StateDB, we propose a new architecture for StateDB, featuring two specialized databases: LiveDB and ArchiveDB.
The two use cases of the databases are summarised in Figure 2(b) and  2(c).
@@ 3. LiveDB with Intrinsic-Pruning
In this section, we describe the design of the LiveDB, which is designed for the validator role. Attributes can be directly changed in their database records without altering a large internal state, as is the case in the MPT data structure (i.e., implies rehashing of changed and new values and their keys for each new or changed attribute). In the LiveDB, hashes are calculated lazily using high-performance computing algorithms such as a lazily executed prefix sum (Ladner and Fischer, 1980) that computes the hash of the worldstate only on demand and reuses hashes of unchanged portions of the worldstate. Attributes have a ’dirty’ flag that indicates whether an attribute has been changed for the lazy hash calculations.
@@
The second design principle is to use binary files that store account and storage attributes in fixed-length records. With a record number, an account attribute, or a storage value can be fetched in constant time by seeking the record and reading the record from the binary file. The seek position is calculated using a linear function over its record number as an argument and the size of the attribute. We have in LiveDB a separate file for each attribute to permit the direct access of attributes via their record number. For the code attribute, we require an extension to fixed-length records, as the code of a smart contract varies and its maximum size is approximately 25 KB.
@@ 4. ArchiveDB for Archival Nodes
In this section, we describe the design of the ArchiveDB, which is specialized for the archive role. It has the same data model as LiveDB, though indexed by a block number that permits the retrieval of a historical block state. The central design concept of ArchiveDB is to store historical state data in sorted tables, building a log of changes over the block numbers for each account and storage keys. Log entries are created if attributes/storage values change. ArchiveDB can employ logs because there are no forks, and for a single block, there can be at most one change of an attribute/storage value. Efficient data structures, such as Log-Structured Merge Trees, are suddenly amenable to forkless chains, although they are not applicable in forked chains.
@@
For the search query, we require additional checks to determine whether the found entry in the log describes the storage location and account of the state query, and whether the account has been deleted before the block. However, the basic design of ArchiveDB is to reduce the lookup of historic information to sorted table lookups. Besides the table for storage values, tables for historical balances, nonces, codes, and account states (both existing and deleted) are required to complete the full historical record. However, these tables have one column less. For example, the table for balance would look like,
@@ 5. Implementation
In this section, we outline the implementation details of our approach. The LiveDB and ArchiveDB have been implemented using Golang. For the LiveDB, we explored three different variants: an in-memory version, a file-based version, and a LevelDB version. Each component and development phase underwent rigorous benchmarking to ensure optimal performance. We used the built-in Golang benchmarking capability or processed sample transactions from the real blockchain. The in-memory database served as our performance upper bound, representing the ideal scenario we aimed to approximate in real-world applications closely. Conversely, the LevelDB database functioned as the performance lower bound, providing a baseline from which we aimed to achieve significantly higher performance compared to current commodity software.
@@ 8. Conclusion
In this paper, we have presented a state database that improves the inefficiency of legacy state databases designed for forking blockchains. Our forkless blockchain database consists of two specialized databases, each tailored to the role of a node in a blockchain. The LiveDB stores data densely and allows overwriting to enable intrinsic pruning, which limits storage overhead. The ArchiveDB is limited to a single subsequent version per block and limits the amount of key copy overhead. Overall, we achieve 100x improvement in storage and a 10x improvement in throughput compared to the geth-based Fantom Blockchain client.

### v2

@@ 1. Introduction
However, forkless blockchains continue to use legacy databases that were initially designed for forking blockchains only. Recent profiling experiences (Baird et al., 2019) report that transaction execution in forkless blockchains is now primarily dominated by storage access, which consumes up to 75% of the block processing. Moreover, they require nodes to store terabytes of data [^1] Storing the entire history of the Ethereum mainnet currently requires 24 TB of disk space, according to https://etherscan.io/chartsync/chainarchive.
@@
Note that the highlighted path for a forkless blockchain is redundant, as it is not needed for the new root $x_{2}$. Retaining the data causes bloat and thus will eventually be pruned (Ethereum, 2025), which typically requires putting the node offline. Moreover, as shown by the dotted arrows, numerous copies of keys along the path need to be copied to create a path from $x_{2}$ to the new value $v^{\prime}_{4}$. Overall, considering an MPT typically holds hundreds of millions of values, the highlighted inefficiencies lead to considerable computational overhead and redundant bloat if previous versions are not needed.
@@
In this paper, we argue that forkless blockchains need forkless blockchain databases. We present a new state database design for forkless blockchains that provides high throughput and low storage costs compared to legacy alternatives.
In our approach, we first specialize the state database according to its intended role in the blockchain. For instance, the LiveDB is a database for validator and observer roles, which require only the state for the latest block on the blockchain. The ArchiveDB is a database for archival roles that keep all historical states for previous blocks.
The key aspect of our database design is that the LiveDB uses a mutable data structure that allows data to be overwritten, reflecting the constrained versioning occurring in forkless blockchains. Overwriting allows us to reflect the state of the most recent execution without keeping stale states in the database (right MPT in Figure 1). Thus, we perform pruning at no additional cost with LiveDB. We refer to this method as intrinsic pruning that is inherently performed as a consequence of the chosen data structure. Moreover, the LiveDB stores key-value maps as densely as possible. A dense representation of the key-value maps results in smaller disk footprints and faster accesses. This design is in stark contrast to the MPT implementations that use key-value stores (e.g., LevelDB (Inc., 2023a), RocksDB (Inc., 2023b)) and SQL databases.
@@
We have implemented LiveDB and ArchiveDB in the Fantom Blockchain and evaluated their utility on real world transactions in a public testnet. Our evaluation shows that LiveDB reduces the storage footprint [^2] Note that even smaller storage decreases imply significant storage reductions linear in the number of nodes in the network because each node replicates the data in a classical L1 blockchain. on the Fantom blockchain by approximately $100\times$ while increasing throughput $10\times$ compared to using the MPT-based Ethereum geth client (Ethereum, 2023).
We summarize our contributions as follows:
@@ 2. StateDB Specialization
The StateDB is necessary to verify the validity of transactions, including those that have been skipped due to insufficient funds, incompatible nonces, and other reasons. Since the validator is tasked only with producing the next block, it does not need to retain any historical state; it only requires the latest state of the StateDB. Therefore, after each processed block of transactions, the state of the StateDB can evolve to the latest state. The observer role is similar to the validator role, but it cannot insert new transactions into the ledger; instead, it participates in the consensus protocol and updates its state accordingly. The archive node role is an extended observer role that keeps the historical states of the blockchain, which can be queried via Remote Procedure Call (RPC). Examples of historical queries include querying an account balance at a specific block height or executing a smart contract at a given block height that queries the information only (but does not change the state). For this reason, it needs to keep the linear evolution of states in the StateDB.
@@
To overcome the shortcomings of this one-size-fits-all StateDB, we propose a new architecture for StateDB, featuring two specialized databases: LiveDB and ArchiveDB. The two use cases of the databases are summarised in Figure 2(b) and  2(c).
@@ 3. LiveDB with Intrinsic-Pruning
In this section, we describe the design of the LiveDB, which is designed for the validator and observer roles.
@@
The second design principle is to use binary files that store account and storage attributes in fixed-length records. With a record number, an account attribute, or a storage value can be fetched in constant time by seeking the record and reading the record from the binary file. The seek position is calculated using a linear function over its record number as an argument and the size of the attribute. In LiveDB there is a separate file for each attribute to permit the direct access of attributes via their record number. For the code attribute, we require an extension to fixed-length records, as the code of a smart contract varies and its maximum size is approximately 25 KB.
@@ 4. ArchiveDB for Archival Nodes
In this section, we describe the design of the ArchiveDB, which is specialized for the archive role.
The ArchiveDB has the same data model as LiveDB, though it is indexed by a block number that permits the retrieval of a historical block state. The central design concept of ArchiveDB is to store historical state data in sorted tables, building a log of changes over the block numbers for each account and storage keys. Log entries are created if attributes/storage values change. ArchiveDB can employ logs because there are no forks, and for a single block, there can be at most one change of an attribute/storage value. Efficient data structures, such as Log-Structured Merge Trees, are suddenly amenable to forkless chains, although they are not applicable in forked chains.
@@
For the search query, we require additional checks to determine whether the found entry in the log describes the storage location and account of the state query, and whether the account has been deleted before the block. However, the basic design of ArchiveDB is to reduce the lookup of historic information to sorted table lookups. Besides the table for storage values, tables for historical balances, nonces, codes, and account states (both existing and deleted) are required to complete the full historical record. However, these tables have one column less. For example, the table for balance would look like:
@@ 5. Implementation
In this section, we outline the implementation details of our approach.
The LiveDB and ArchiveDB have been implemented using Golang. For the LiveDB, we explored three different variants: an in-memory version, a file-based version, and a LevelDB version. Each component and development phase underwent rigorous benchmarking to ensure optimal performance. We used the built-in Golang benchmarking capability or processed sample transactions from the real blockchain. The in-memory database served as our performance upper bound, representing the ideal scenario we aimed to approximate in real-world applications closely. Conversely, the LevelDB database functioned as the performance lower bound, providing a baseline from which we aimed to achieve significantly higher performance compared to current commodity software.
@@ 8. Conclusion
In this paper, we have presented a state database specialized for forkless blockchains. Compared to legacy state databases our solution provides improvements both in terms of throughput as well as storage overhead. Our forkless blockchain database consists of two specialized databases, each tailored to the role of a node in a blockchain. The LiveDB stores data densely and allows overwriting to enable intrinsic pruning, which limits storage overhead. The ArchiveDB is limited to a single subsequent version per block and limits the amount of key copy overhead. Overall, we report a 100x improvement in storage and a 10x improvement in throughput compared to the geth-based Fantom Blockchain client.
9. Acknowledgment
We would like to thanks following colleagues and engineers for helping with implementing this work, and for their valuable feedback: Jiri Malek, Jan Kalina, Lukas Pelanek.